he lengths of the three sides of a triangle (not necessarily a right triangle) are 3.56 meters, 5.07 meters and 2.09 meters. What is the cosine of the angle opposite the side of length 2.09 meters?

cos theta =_________
Answer #1

cos theta = b/c = 3.56/5.07 = 0.702

so, theta = 45.39 degrees